# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ 3.1. Raw events
Individual Events. Raw events $e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})$ are utilized by event-by-event processing methods such as probabilistic filters and Spiking Neural Networks (SNNs). These methods build additional information from past events or external knowledge and fuse it with incoming events asynchronously to produce an output [72, 73].
Event Packet. The event set $E=\{e_{k}\}^{N_{e}}$ retains precise timestamp and polarity information for each event. Selecting the appropriate packet size $N_{e}$ is crucial to meet the assumptions of the algorithm (e.g., constant motion speed throughout packet’s duration), which varies depending on the task [74, 75]
@@ 3.2. Event frame (2D Grid)
Histogram. This representation converts events into histograms, offering an activity-driven sample rate. Although not fully aligned with the event-based paradigm, it has proven effective [76, 77]. Traditional 2D histograms discretize events into bins, while the Activity-Aware Event Integration Module extends them with spatiotemporal operations, capturing finer details and improving dynamic-scene performance, particularly in semantic segmentation [78].
Time Surface. A Time Surface (TS) is a 2D map where each pixel stores a single time value (e.g., the timestamp of the last event at that pixel). Events are converted into an image whose "intensity" is a function of the motion history at that location, with larger values corresponding to more recent motion. TSs explicitly expose the rich temporal information of the events and can be updated asynchronously [79, 80, 81].
Figure 5. Event representations method (a) Raw event (b) 2D histogram (c) Time surface (d) Voxel grid (e) RGB picture.
@@ 3.3. Spatio-temporal 3D grid representation
A Voxel Grid is a 3D space-time histogram where each voxel corresponds to a pixel and time interval, preserving temporal information better than 2D projections. With polarity, it forms a discretized scalar field on the image plane, using accumulation or kernel distribution for sub-voxel accuracy [82, 83, 84]. This representation retains rich spatiotemporal detail for precise event tracking but requires more memory and computation, posing challenges for real-time use.
@@ 3.4. Customized representation
Customized event representations address task-specific limitations of traditional methods by combining spatial, temporal, or domain-specific features to boost efficiency and accuracy. For example, [85] employs adaptive filtering for motion deblurring to preserve high-frequency details, while [86] introduces the 2D-1T Event Cloud Sequence (ECS), which separates spatial and temporal components to retain sparsity and capture both geometry and motion, aiding recognition.
Task-specific designs improve performance by exploiting relevant features and domain knowledge, but their generalizability across tasks remains limited. For example, the adaptive filtering representation in [85] is tightly coupled with a generative model and an Extended Kalman Filter, making it powerful for dynamic tracking but hard to adapt to recognition or reconstruction tasks that require spatially structured outputs. Similarly, the ECS representation [86] yields compact embeddings suitable for classification but lacks the explicit geometric and kinematic detail needed for precise tracking or mapping. Moreover, both are deeply embedded in their respective architectures, so transferring them to other paradigms demands significant redesign and retraining. Consequently, while customized representations demonstrate how tailoring to specific assumptions can achieve state-of-the-art performance, they also reveal the trade-off between specialization and generalization. Future research toward more unified or transferable event representations, along with standardized benchmarks for cross-task robustness, would be valuable for broadening applicability.
@@ 4.1. Event-based denoising
Literature review. Event denoising has achieved notable advancements through diverse methodologies, including statistical approaches, filtering-based techniques, surface fitting and deep learning, significantly improving noise suppression in event-based vision (Fig. 6). To ensure fair comparison among these methods despite their diverse assumptions, we adopt the large-scale real-world LED dataset [94], which provides paired noisy/clean events under controllable illumination and noise levels. All methods are evaluated using the same metric denoising accuracy $DA=\frac{1}{2}(\frac{TP}{GP}+\frac{TN}{GN})$, which decomposes into signal retention (SR) and noise removal (NR), thus jointly capturing the two key aspects of denoising quality. As illustrated in Fig. 6, this unified setting enables an objective and reproducible comparison of statistical, filtering-based, and deep learning strategies.
@@
Statistical methods. Early techniques relied on statistical analysis to identify outliers by evaluating event density in local spatio-temporal neighborhoods [95]. Delbruck et al. [96] pioneered density-based filtering, leveraging local context to suppress noisy events. Subsequent efforts [89, 90] enhanced these techniques with optimized event storage strategies to reduce computational complexity and improve processing efficiency. However, these approaches often require manual parameter tuning to adapt to varying noise conditions, limiting their scalability and generalizability.
@@
Filtering-based methods. To better exploit the temporal and asynchronous nature of event data, several filtering algorithms have been developed. (i) Temporal filters [97] remove redundant events by leveraging temporal correlations. (ii) Spatial filters [98] isolate motion-related events by analyzing pixel intensity changes. (iii) Spatio-temporal filters [95, 99] combine both strategies to suppress background activity (BA) noise. For example, Liu et al. [100] showed that integrating spatial and temporal filtering reduces BA noise while preserving critical motion events.
@@
Surface fitting techniques. Surface fitting methods offer an alternative approach, particularly effective for smoothing event data in continuous motion scenarios. Methods like EV-Gait [91] and the Guided Event Filter (GEF) [101] employ local plane fitting , optical flow estimation, and image gradients for noise smoothing. The time surface (TS) method [88, 97] transforming event streams into monotonically decreasing representations, effectively mitigates sparsity but exhibits limited performance in low light or highly dynamic scenes.
@@
Deep learning-based methods. Recent advances in deep learning have revolutionized event denoising, enabling automated solutions by training on noisy-clean event pairs [102, 103]. Sparse feature learning (K-SVD [104], MLPF [90]) methods focus on sparse feature extraction and event probability estimation. EDnCNN [92] integrates frame and IMU data to classify events as signal or noise. EventZoom [105] employs a U-Net architecture for efficient noise-to-noise denoising, achieving superior performance in handling noisy event streams. AEDNet [93] processes raw DVS data while preserving inherent spatio-temporal correlations. EDformer [106] introduces an event-by-event transformer model, while EDmamba [107] leverages state space modeling to achieve efficient noise-aware spatiotemporal denoising.
@@ 4.2. Event-based filtering and feature extraction
Motivation. Unlike traditional frame-based cameras capturing images at fixed intervals, event cameras operate asynchronously, producing sparse data only when significant brightness changes occur. This dynamic output demands specialized filtering to isolate meaningful events for effective feature extraction. Event filtering highlights important scene dynamics—like motion and edges—while removing irrelevant or redundant data. This enhances efficiency and accuracy in tasks such as object detection, motion tracking, and scene reconstruction, especially in fast-moving or rapidly changing lighting conditions. By reducing noise and preserving critical features, filtering supports robust, low-latency processing essential for real-time applications.
Challenge. The asynchronous nature of event data poses key challenges for filtering, as there’s no universal definition of a “significant” event. Variations in event frequency and timing complicate maintaining temporal coherence, which is vital for accurate motion representation during feature extraction. Balancing filtering complexity with real-time processing is also difficult—advanced methods may boost accuracy but risk adding latency. Parameter tuning is another challenge, since optimal settings vary with environmental conditions and must adapt to dynamic, unpredictable scenes. This need for adaptability makes robust filtering crucial for reliable event-based vision. Overcoming these issues is essential to enhance feature extraction and ensure consistent real-world performance.
@@
Frame-based methods [108, 109]
@@
Loss of fine event details
@@
Initial event processing
Surface of Active Events-based methods [110, 111]
@@
Good feature localization
@@
Complex computational requirements
@@
Clustering-based methods [112, 113]
@@
Fast stream processing
@@
Dynamic scene analysis
Geometric transform-based methods [114, 115, 116]
@@
Robust feature detection
@@
High resource consumption
@@
Extreme illumination scenarios
Temporal filtering-based methods [117, 118]
@@
Good real-time performance
@@
Detail preservation issues
@@
Resource-constrained systems
Asynchronous methods [119, 120, 121]
@@
Efficient event handling
@@
Noise-heavy environments
Hybrid methods [122, 123, 124]
@@
Multi-layer filtering
@@
Complex parameter tuning
@@
Multi-feature scenarios
Neural networks-based methods [125, 126, 117, 127]
@@
Accuracy-speed tradeoffs
@@
Biological vision systems
Frame-event hybrid methods [128, 129, 130]
@@
Complex resource management
@@
High-speed tracking
@@
Literature review. Event-based feature extraction has progressed from adapting frame-based methods to developing specialized event-driven approaches (Tab. 6). Early work modified traditional techniques, such as applying Harris corner detection [108] to event accumulation frames. A major advance was the Surface of Active Events (SAE) [110], which records the timestamp of the latest event per pixel, preserving temporal precision while improving efficiency and feature extraction performance.
@@
Feature extraction methods have advanced in both accuracy and efficiency. eFAST [119] shifted from gradient-based to faster comparison-based operations tailored for event data, while Arc* [111] improved detection speed and repeatability through refined SAE filtering. Hybrid methods like FA-Harris [122] and TLF-Harris [123] balanced efficiency and robustness via candidate selection and multi-layer filtering. FEAST [125] introduced unsupervised extraction with spiking neuron-like units, and ROT-Harris [124] used tree-based processing to surpass traditional 2D approaches. Together, these works highlight a shift toward practical, scalable implementations.
@@
Parallel to advancements in feature extraction, event filtering has emerged as a critical preprocessing step for robust feature extraction, leveraging the asynchronous and high-resolution temporal data of event cameras. Temporal filtering-based methods have significantly improved data quality and computational efficiency by prioritizing meaningful scene changes [117, 118]. Clustering-based methods, such as eCDT [112], dynamically group events to represent dense streams compactly while minimizing complexity. However, such methods often face challenges in maintaining robustness under varying conditions due to sensitivity to parameter configurations [113].
@@
Parametric filtering has further advanced event-based vision by employing geometric transformations to filter irrelevant data. For instance, EKLT [114] aligns events to improve feature extraction, while curve-fitting techniques construct smooth spatio-temporal trajectories [115], excelling in scenarios with rapid motion or extreme illumination. Asynchronous methods, such as HASTE [120], process individual events in real time using hypothesis-driven transformations, and proximity-based trackers continuously refine feature locations, effectively suppressing noise and supporting low-latency applications [121].
@@
Integrating filtering with feature extraction has greatly improved the robustness and efficiency of event-based vision systems. For instance, combining filters with classical algorithms like Harris or FAST enables reliable corner detection, while shape detection methods using ICP and Hough transforms enhance performance in high-speed, high-dynamic-range settings [128]. Modern tracking approaches also exploit the low latency and high temporal resolution of event cameras by incorporating probabilistic associations and spatio-temporal constraints, resulting in more stable features and improved tracking accuracy [129, 130].
@@
Neural network-based filtering integrates asynchronous event processing with biologically inspired architectures. Pulse-based neural networks [126] exploit the high temporal resolution of event data for precise feature tracking, showing strong potential in dynamic environments. Yet, current models still fall short of the detection accuracy achieved by convolutional networks [117, 127]. Ongoing efforts to unify filtering, feature extraction, and tracking are pushing event-based vision toward broader applicability while preserving efficiency and robustness in real-world scenarios.

### v4

@@ 3.1. Raw events
Individual Events. Raw events $e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})$ are utilized by event-by-event processing methods such as probabilistic filters and Spiking Neural Networks (SNNs). These methods build additional information from past events or external knowledge and fuse it with incoming events asynchronously to produce an output [wang2025eas, guo2025event].
Event Packet. The event set $E=\{e_{k}\}^{N_{e}}$ retains precise timestamp and polarity information for each event. Selecting the appropriate packet size $N_{e}$ is crucial to meet the assumptions of the algorithm (e.g., constant motion speed throughout packet’s duration), which varies depending on the task [lin2024compressed, li2024gaze]
@@ 3.2. Event frame (2D Grid)
Histogram. This representation converts events into histograms, offering an activity-driven sample rate. Although not fully aligned with the event-based paradigm, it has proven effective [tang2025spike, kim2022real]. Traditional 2D histograms discretize events into bins, while the Activity-Aware Event Integration Module extends them with spatiotemporal operations, capturing finer details and improving dynamic-scene performance, particularly in semantic segmentation [10477577].
Time Surface. A Time Surface (TS) is a 2D map where each pixel stores a single time value (e.g., the timestamp of the last event at that pixel). Events are converted into an image whose "intensity" is a function of the motion history at that location, with larger values corresponding to more recent motion. TSs explicitly expose the rich temporal information of the events and can be updated asynchronously [10401990, grimaldi2024robust, huang2023progressive].
Figure 5. Event representations method. (a) Raw event. (b) 2D histogram. (c) Time surface. (d) Voxel grid. (e) RGB picture.
@@ 3.3. Spatio-temporal 3D grid representation
A Voxel Grid is a 3D space-time histogram where each voxel corresponds to a pixel and time interval, preserving temporal information better than 2D projections. With polarity, it forms a discretized scalar field on the image plane, using accumulation or kernel distribution for sub-voxel accuracy [gehrig2024dense, cao2025embracing, xu2024dmr]. This representation retains rich spatiotemporal detail for precise event tracking but requires more memory and computation, posing challenges for real-time use.
@@ 3.4. Customized representation
Customized event representations address task-specific limitations of traditional methods by combining spatial, temporal, or domain-specific features to boost efficiency and accuracy. For example, [wang2024asynchronous] employs adaptive filtering for motion deblurring to preserve high-frequency details, while [chen2022ecsnet] introduces the 2D-1T Event Cloud Sequence (ECS), which separates spatial and temporal components to retain sparsity and capture both geometry and motion, aiding recognition.
Task-specific designs improve performance by exploiting relevant features and domain knowledge, but their generalizability across tasks remains limited. For example, the adaptive filtering representation in [wang2024asynchronous] is tightly coupled with a generative model and an extended kalman filter, making it powerful for dynamic tracking but hard to adapt to recognition or reconstruction tasks that require spatially structured outputs. Similarly, the ECS representation [chen2022ecsnet] yields compact embeddings suitable for classification but lacks the explicit geometric and kinematic detail needed for precise tracking or mapping. Moreover, both are deeply embedded in their respective architectures, so transferring them to other paradigms demands significant redesign and retraining. Consequently, while customized representations demonstrate how tailoring to specific assumptions can achieve state-of-the-art performance, they also reveal the trade-off between specialization and generalization. Future research toward more unified or transferable event representations, along with standardized benchmarks for cross-task robustness, would be valuable for broadening applicability.
@@ 4.1. Event-based denoising
Literature review. Event denoising has achieved notable advancements through diverse methodologies, including statistical approaches, filtering-based techniques, surface fitting and deep learning, significantly improving noise suppression in event-based vision (Fig. 6). To ensure fair comparison among these methods despite their diverse assumptions, we adopt the large-scale real-world LED dataset [duan2024led], which provides paired noisy/clean events under controllable illumination and noise levels. All methods are evaluated using the same metric denoising accuracy $DA=\frac{1}{2}(\frac{TP}{GP}+\frac{TN}{GN})$, which decomposes into signal retention (SR) and noise removal (NR), thus jointly capturing the two key aspects of denoising quality. As illustrated in Fig. 6, this unified setting enables an objective and reproducible comparison of statistical, filtering-based, and deep learning strategies.
@@
Statistical methods. Early techniques relied on statistical analysis to identify outliers by evaluating event density in local spatio-temporal neighborhoods [feng2020event]. Delbruck et al. [delbruck2008frame] pioneered density-based filtering, leveraging local context to suppress noisy events. Subsequent efforts [khodamoradi2018n, guo2022low] enhanced these techniques with optimized event storage strategies to reduce computational complexity and improve processing efficiency. However, these approaches often require manual parameter tuning to adapt to varying noise conditions, limiting their scalability and generalizability.
@@
Filtering-based methods. To better exploit the temporal and asynchronous nature of event data, several filtering algorithms have been developed. (i) Temporal filters [baldwin2019inceptive] remove redundant events by leveraging temporal correlations. (ii) Spatial filters [ieng2014asynchronous] isolate motion-related events by analyzing pixel intensity changes. (iii) Spatio-temporal filters [feng2020event, guo2021hashheat] combine both strategies to suppress background activity (BA) noise. For example, Liu et al. [liu2015design] showed that integrating spatial and temporal filtering reduces BA noise while preserving critical motion events.
@@
Surface fitting techniques. Surface fitting methods offer an alternative approach, particularly effective for smoothing event data in continuous motion scenarios. Methods like EV-Gait [wang2019ev] and the Guided Event Filter (GEF) [duan2021guided] employ local plane fitting , optical flow estimation, and image gradients for noise smoothing. The time surface (TS) method [lagorce2016hots, baldwin2019inceptive] transforming event streams into monotonically decreasing representations, effectively mitigates sparsity but exhibits limited performance in low light or highly dynamic scenes.
@@
Deep learning-based methods. Recent advances in deep learning have revolutionized event denoising, enabling automated solutions by training on noisy-clean event pairs [ruan2025pre, ruan2024distill]. Sparse feature learning (K-SVD [xie2018dvs], MLPF [guo2022low]) methods focus on sparse feature extraction and event probability estimation. EDnCNN [baldwin2020event] integrates frame and IMU data to classify events as signal or noise. EventZoom [duan2021eventzoom] employs a U-Net architecture for efficient noise-to-noise denoising, achieving superior performance in handling noisy event streams. AEDNet [fang2022aednet] processes raw DVS data while preserving inherent spatio-temporal correlations. EDformer [jiang2024edformer] introduces an event-by-event transformer model, while EDmamba [ruan2025edmamba] leverages state space modeling to achieve efficient noise-aware spatiotemporal denoising.
@@ 4.2. Event-based filtering and feature extraction
Motivation. Unlike traditional frame-based cameras capturing images at fixed intervals, event cameras operate asynchronously, producing sparse data only when brightness changes occur. This dynamic output demands specialized filtering to isolate meaningful events for effective feature extraction. Event filtering highlights important scene dynamics (e.g., motion and edges) while removing irrelevant or redundant data. This enhances efficiency and accuracy in tasks such as object detection, motion tracking, and scene reconstruction, especially in fast-moving or rapidly changing lighting conditions. By reducing noise and preserving critical features, filtering supports robust, low-latency processing essential for real-time applications.
Challenge. The asynchronous nature of event data poses key challenges for filtering, as there’s no universal definition of a “significant” event. Variations in event frequency and timing complicate maintaining temporal coherence, which is vital for accurate motion representation during feature extraction. Balancing filtering complexity with real-time processing is also difficult, advanced methods may boost accuracy but risk adding latency. Parameter tuning is another challenge, since optimal settings vary with environmental conditions and must adapt to dynamic, unpredictable scenes. This need for adaptability makes robust filtering crucial for reliable event-based vision. Overcoming these issues is essential to enhance feature extraction and ensure consistent real-world performance.
@@
Frame-based methods [harris1988combined, vasco2016fast]
@@
Loss of fine event details.
@@
Initial event processing.
Surface of Active Events-based methods [benosman2013event, 8491018]
@@
Good feature localization.
@@
Complex computational requirements.
@@
Clustering-based methods [hu2022ecdt, gao2024sd2event]
@@
Fast stream processing.
@@
Dynamic scene analysis.
Geometric transform-based methods [gehrig2020eklt, seok2020robust, chui2021event]
@@
Robust feature detection.
@@
High resource consumption.
@@
Extreme illumination scenarios.
Temporal filtering-based methods [yi2023deep, wan2024event]
@@
Good real-time performance.
@@
Detail preservation issues.
@@
Resource-constrained systems.
Asynchronous methods [mueggler2017fast, alzugaray2020haste, chiberre2022long]
@@
Efficient event handling.
@@
Noise-heavy environments.
Hybrid methods [li2019FAHarris, mohamed2021dynamic, harrigan2021rot]
@@
Multi-layer filtering.
@@
Complex parameter tuning.
@@
Multi-feature scenarios.
Neural networks-based methods [afshar2020event, barchid2023spiking, yi2023deep, yao2023attention]
@@
Accuracy-speed tradeoffs.
@@
Biological vision systems.
Frame-event hybrid methods [best1992method, zhu2017event, rodriguez2020asynchronous]
@@
Complex resource management.
@@
High-speed tracking.
@@
Literature review. Event-based feature extraction has progressed from adapting frame-based methods to developing specialized event-driven approaches (Tab. 6). Early work modified traditional techniques, such as applying Harris corner detection [harris1988combined] to event accumulation frames. A major advance was the Surface of Active Events (SAE) [benosman2013event], which records the timestamp of the latest event per pixel, preserving temporal precision while improving efficiency and feature extraction performance.
@@
Feature extraction methods have advanced in both accuracy and efficiency. eFAST [mueggler2017fast] shifted from gradient-based to faster comparison-based operations tailored for event data, while Arc* [8491018] improved detection speed and repeatability through refined SAE filtering. Hybrid methods like FA-Harris [li2019FAHarris] and TLF-Harris [mohamed2021dynamic] balanced efficiency and robustness via candidate selection and multi-layer filtering. FEAST [afshar2020event] introduced unsupervised extraction with spiking neuron-like units, and ROT-Harris [harrigan2021rot] used tree-based processing to surpass traditional 2D approaches. Together, these works highlight a shift toward practical, scalable implementations.
@@
Parallel to advancements in feature extraction, event filtering has emerged as a critical preprocessing step for robust feature extraction, leveraging the asynchronous and high-resolution temporal data of event cameras. Temporal filtering-based methods have significantly improved data quality and computational efficiency by prioritizing meaningful scene changes [yi2023deep, wan2024event]. Clustering-based methods, such as eCDT [hu2022ecdt], dynamically group events to represent dense streams compactly while minimizing complexity. However, such methods often face challenges in maintaining robustness under varying conditions due to sensitivity to parameter configurations [gao2024sd2event].
@@
Parametric filtering has further advanced event-based vision by employing geometric transformations to filter irrelevant data. For instance, EKLT [gehrig2020eklt] aligns events to improve feature extraction, while curve-fitting techniques construct smooth spatio-temporal trajectories [seok2020robust], excelling in scenarios with rapid motion or extreme illumination. Asynchronous methods, such as HASTE [alzugaray2020haste], process individual events in real time using hypothesis-driven transformations, and proximity-based trackers continuously refine feature locations, effectively suppressing noise and supporting low-latency applications [chiberre2022long].
@@
Integrating filtering with feature extraction has greatly improved the robustness and efficiency of event-based vision systems. For instance, combining filters with classical algorithms like Harris or FAST enables reliable corner detection, while shape detection methods using ICP and Hough transforms enhance performance in high-speed, high-dynamic-range settings [best1992method]. Modern tracking approaches also exploit the low latency and high temporal resolution of event cameras by incorporating probabilistic associations and spatio-temporal constraints, resulting in more stable features and improved tracking accuracy [zhu2017event, rodriguez2020asynchronous].
@@
Neural network-based filtering integrates asynchronous event processing with biologically inspired architectures. Pulse-based neural networks [barchid2023spiking] exploit the high temporal resolution of event data for precise feature tracking, showing strong potential in dynamic environments. Yet, current models still fall short of the detection accuracy achieved by convolutional networks [yi2023deep, yao2023attention]. Ongoing efforts to unify filtering, feature extraction, and tracking are pushing event-based vision toward broader applicability while preserving efficiency and robustness in real-world scenarios.